OpenStudy (anonymous):

Find four consecutive integers whose sum is 114.

OpenStudy (anonymous):

The first integer let's call A The one after that is A+1, then A+2, then A+3... with me so far?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

Therefore A + A + 1 + A + 2 + A + 3 = 114 can you go from there?

OpenStudy (anonymous):

I can try yes

OpenStudy (anonymous):

okay tell me what you get.

OpenStudy (anonymous):

well I added 1+2+3

OpenStudy (anonymous):

and got 6

OpenStudy (anonymous):

okay good. when you add the A's together you should get 4A + 6 = 114. Solve A.

OpenStudy (anonymous):

here... \[A + A + 1 + A + 2 + A + 3 = 114\] \[4A = 108\]\[A = 108\div4 = 27\] A2 = A + 1. A3 = A + 2 A4 = A+3 so what is your final answer?